the circumference of a rectanglen is 70.if lenth is 7 times of height then wt is the area of rectangle
Answer by Alan3354(69443)

the circumference of a rectanglen is 70.if lenth is 7 times of height then wt is the area of rectangle
----
C = 2L + 2H = 70
L + H = 35
L = 7H
-------
Sub for L in L + H = 35
7H + H = 35
H = 35/8
L = 245/8
---
Area = L*H = 8575/64 sq units